OpenStudy (anonymous):

Which one is the biggest between .1 and .11111...

OpenStudy (anonymous):

.11111...

OpenStudy (anonymous):

.1111.... can be rounded of to just .1 but technically .1111... is bigger

OpenStudy (anonymous):

for solving such questions, we first make the number of digits after the decimal point same by just adding zeros as follows : 0.10000 and 0.11111 now convert them to decimal numbers as follows 10000 11111 -------- and --------- 100000 100000 denominator is same now, so we just compare the numerators and find 10000 < 11111 so 10000 11111 -------- < --------- 100000 100000 Hence 0.1 < 0.11111

OpenStudy (anonymous):

Thank you guys. @Harkirat - Nice description :)

OpenStudy (anonymous):

u r welcome....☺

OpenStudy (jamesj):

Alternatively \[0.1 = \frac{1}{10}\] and\[0.1111... = \frac{1}{10} +\frac{1}{100} +\frac{1}{1000} +\frac{1}{10000} + ...\] Since all the additional terms after the 1/10 are also positive, it must be the case that \[0.1111... > 0.1\]

OpenStudy (jamesj):

Alternatively,\[0.111... - 0.1 = 0.0111... > 0\] hence \[0.111... > 0.1\] So as you can see, there's lots of ways to attack this problem.

OpenStudy (anonymous):

@ JamesJ Yes, your answer also r great !!!! Medal for u ...☺